# Gen-C: Populating Virtual Worlds with Generative Crowds

### v3

@@ Abstract
Over the past two decades, researchers have made significant steps in simulating agent-based human crowds, yet most efforts remain focused on low-level tasks such as collision avoidance, path following, and flocking. Realistic simulations, however, require modeling high-level behaviors that emerge from agents interacting with each other and with their environment over time. We introduce Generative Crowds (Gen-C), a generative framework that produces crowd scenarios capturing agent–agent and agent–environment interactions, shaping coherent high-level crowd plans. To avoid the labor-intensive process of collecting and annotating real crowd video data, we leverage large language models (LLMs) to bootstrap synthetic datasets of crowd scenarios. We propose a time-expanded graph representation, encoding actions, interactions, and spatial context. Gen-C employs a dual Variational Graph Autoencoder (VGAE) architecture that jointly learns connectivity patterns and node features conditioned on textual and structural signals, overcoming the limitations of direct LLM generation to enable scalable, environment-aware multi-agent crowd simulations. We demonstrate the effectiveness of Gen-C on scenarios with diverse behaviors such as, a University Campus and a Train Station, showing that it generates heterogeneous crowds, coherent interactions, and high-level decision-making patterns consistent with real-world crowd dynamics.
@@ 1 Introduction
Simulating human crowds has been an active area of research for several years, leading to a wide range of models and tools for generating realistic human behaviors [1, 2]. Several commercial solutions have also been integrated into game engines and rendering software to support large-scale crowd scenes [3, 4, 5, 6]. These approaches have advanced the realism of simulation, but they primarily target motion synthesis rather than the decision processes and interactions that govern behaviors over time.
At the algorithmic level, most methods have concentrated on low-level navigation tasks such as collision avoidance, path following, and group steering [7, 8]. While such methods effectively handle local interactions, they fall short of capturing the complex behaviors that arise when individuals pursue goals, interact socially, and respond to the affordances of their environment. For example, people stop to talk with friends, browse shop windows, or queue to buy tickets before boarding a train. Such behaviors need planning and coordination and are rarely modeled in current systems, which leads to repetitive dynamics that lack realism and adaptability.
@@
In this work, we introduce Generative Crowds (Gen-C), a framework to automatically generate high-level behaviors of virtual agents, enabling multi-character simulations with diverse crowds. Throughout this paper, we use the term agent to denote a virtual character navigating within the crowd. At the core of our approach lies a crowd scenario graph: a time-expanded graph encoding both spatial and temporal evolution of the crowd, capturing high-level agent-to-agent and agent-to-environment interactions over time. With the advent of big data and deep learning, an obvious solution would be to learn such crowd scenario graphs from real data. Several data-driven methods have indeed modeled crowd by leveraging trajectories extracted from limited real-world datasets [9, 10, 11, 12, 13]. Nevertheless, the challenges and dependency on collecting and annotating data can quickly become a barrier and also limits the generalization capabilities of such solutions.
To break free from this dependency, we propose to automatically synthesize crowd datasets, featuring high-level behaviors, by leveraging the power of Large Language Models (LLMs) like the GPT framework [14]. While LLMs can in principle be used directly to script crowd scenarios, as we will show, this approach faces three main drawbacks: it scales poorly beyond a few agents, leading to repetitive behavior; it depends on heavy prompt engineering and domain expertise to produce coordinated behaviors; and it lacks efficiency for generating structured responses, requiring access to large and costly models. We therefore use LLMs only to create a small seed set of representative crowd scenarios that describe agent-agent and agent-environment interactions for given input scenes, which we subsequently represent as crowd scenario graphs.
Given such graphs, a generative model is trained to learn a distribution of crowd behaviors and interactions that we can then sample and synthesize new graphs. As most existing graph embedding algorithms are mainly developed for static graphs, they cannot capture the complex and dynamic nature of a crowd scenario graph. To that end, we propose a model architecture that leverages two Variational Graph Autoencoders (VGAEs) [15] to capture both how agents interact with each other and with their environment. The trained model can be conditioned on natural language, allowing users to populate scenes with structured crowds simulations directly from text.
@@
Our approach builds upon recent advances in generative AI, where models have been successfully applied to text-driven human motion synthesis [16] and agent or robot planning [17]. However, these methods mainly focus on single-agent scenarios [18, 19, 20, 21, 22]. Closer to our work, text-based diffusion models have been explored for coordinating groups [23] and synthesizing crowd trajectories [24], but these methods do not capture structured, high-level behaviors. In contrast, Gen-C serves as a generative framework for high-level behavioral planning among multiple agents in a crowd, generating structured interaction graphs that model both agent–agent and agent–environment interactions, with a design extensible to text-guided generation. Overall, Gen-C introduces:
  1. 1.
    A graph-based representation that captures temporal agent-agent and agent-environment interactions.
  2. 2.
    A text-conditioned dual VGAE that jointly learns graph structure and features, enabling scalable generation of novel multi-agent scenarios.
  3. 3.
    A synthetic data pipeline, leveraging LLMs to bootstrap crowd scenarios, reducing reliance on real annotations.
@@ 2 Related Work
Crowd Simulation. Crowd simulation has a long history, with early works introducing rule-based models for navigation and flocking [25, 26]. Subsequent research proposed higher-level control mechanisms, such as crowd patches and sketch-based navigation fields [27, 28, 29], which allowed users to shape collective motion in structured environments. A major thrust of later work has been to increase realism by modeling heterogeneity, for example through group-level properties [30], individual personality traits [31, 32, 33], or stylistic variations across different types of crowds [34]. These methods highlighted the importance of diversity for believable simulations but typically required manual specification or extensive parameter tuning. In parallel, data-driven approaches emerged that learn behavioral patterns from real-world crowd trajectories [35, 36, 37], sometimes adopting graph-based representations [38, 39] that connect crowd modeling to broader developments in machine learning. More recently, reinforcement learning and hybrid imitation-learning methods [40, 13, 24, 12, 41] and diffusion-based generative approaches [42, 43] have demonstrated the potential of scalable and adaptive policies that capture richer dynamics. Despite these advances, most existing methods concentrate on low-level navigation or local diversity. They do not explicitly model structured, high-level multi-agent behaviors that depend on both agent-agent and agent-environment interactions over extended horizons. Our work addresses this gap by introducing a framework for generating heterogeneous, scenario-driven crowd behaviors that capture such multi-layered interactions.
@@
Large Language Models and Autonomous Agents. LLMs have shown to be effective to domains beyond Natural Language Processing (NLP) such as task planning, embodied reasoning, autonomous agents, and floor plan design [44, 45, 46, 47, 48, 49]. Here, we explore an architecture that allows a LLM to generate a sequence of crowd behaviors, over time and space, that forms synthetic crowd datasets. We further exploit the power of generative NLP models to enable text-guided synthesis of high-level crowd behaviors. Along these lines a lot of work has explored generative models for text-driven human motion generation tasks [16]. Despite impressive results, including generating scene-and-language conditioned motions [18, 19, 20, 21, 22], such methods typically focus on animating individual characters. But a human crowd is much more than that, consisting of multiple agents performing a sequence of actions while interacting with other agents and the environment. Along these lines, the recent work of Ji et al. [23] uses a text-based diffusion model, in combination with LLMs, to steer crowds of agents but does not extend to other types of behaviors. Our work seeks to learn high-level behaviors captured via crowd scenario graphs that can complement the low-level steering capabilities of such agents.
@@
Graph Generation. Beyond our domain, learning graph representations is a fundamental task across many real-world applications [50]. While much prior work has focused on learning static graphs using deep learning models such as VGAEs [15], recent research has extended these ideas to dynamic graphs, where the structure evolves over time. Approaches include neural temporal point processes and attention mechanisms for modeling time-dependent events [51], as well as deep architectures such as Temporal Graph Networks [52] and Spatio-Temporal Graph Convolutional Networks [53] for capturing temporal dependencies and predicting future states. While related, our work focuses on crowd scenario graphs: dynamic graphs with rich relational features that capture the evolution of crowds. Unlike most dynamic graph models aimed at tasks such as link prediction, node classification, or recommendation, our formulation targets the evolution of crowd interactions. Closely related, conditional graph generation has also gained attention, with most approaches generating only the graph structure while assuming fixed node and edge features. For instance, NGG [54] employs conditioned latent diffusion, GraphVAE [55] integrates conditional codes for guided generation, and MOOD [56] guides diffusion sampling with property gradients. Sequential models [57] use conditional inputs to initialize node and edge creation, while others embed conditional codes within message-passing schemes [58]. Complementary to these efforts, the Gen-C framework models both graph structures and node features via distinct, condition-aware latent distributions, enabling generation directly from textual input.
@@ 3 Overview
Preliminaries. A crowd scenario graph represents a crowd scenario, consisting of agents, a series of events, environment locations, and the associated actions and interactions. To simplify the generation and learning process, we define a finite set of actions and locations which, when combined, can describe a wide range of crowd behaviors. Specifically, an Action is a label from $\mathbf{\mathcal{A}ct}=$~{wait, sit, wander, queue, enter/exit, wave at, discuss, meet, object interact, talk on phone, read, look at, carry} that specifies an agent’s behavior at a given time. When multiple agents perform a shared action (e.g., discuss), this is referred to as an Interaction. An Environment Location is defined by its name, category, position, and scale. While the name may be abstract, the category is selected from a finite set: $\mathbf{\mathcal{L}oc}=$~{building, room, entrance, exhibit, furniture, outdoor area, item, service area}. Finally, an Event encodes a sequence of actions together with their location and participating agents. Specifically, each event consists of a set of action–location pairs (e.g., {discuss, room}), the participating agents, and a trigger condition that determines its initiation (i.e., starts after a specific event).
The Gen-C Framework. Our framework for generating crowd scenarios consists of three main components: Synthetic Data Generation, Learning, and Scenario Generation. First, we use a LLM to produce diverse synthetic scenarios through targeted queries that contains events describing agent actions, interactions, and environment placements. These scenarios are transformed into graph structures encoding action sequences, relationships, and interactions. Second, we train two synergistic VGAE models: one reconstructs the graph structure (agent interactions), while the other defines node features (agent actions and navigation). A text-conditioned prior integrates natural language, aligning the generated graphs with input descriptions. Finally, novel scenarios are synthesized by sampling from the learned latent distributions, producing contextually consistent multi-agent behaviors.
@@ 4 Synthetic Data Generation
Simulating diverse high-level crowd behaviors requires extensive high-quality data, yet collecting and annotating real crowds is costly and labor-intensive. To address this, we leverage recent advances in LLMs to synthesize data consistent with crowd dynamics.
@@ 4.1 Bootstrapping Crowd Scenarios via LLMs
To generate simulations that capture representative crowd dynamics, we design two tailored LLM queries, each serving a specific purpose. The process begins with a single-sentence input $S_{in}$ describing a crowd scenario, which serves as the seed for generation. We construct a set of $100$ such seeds and, for each, create $50$ paraphrased variants to increase diversity, yielding $5000$ textual inputs in total per theme; in this work we introduce two themes, University Campus and Train Station. Each sentence encodes three main components—a location phrase, one or more subject noun phrases, and one or more action verb phrases—without following a fixed syntactic template. We avoid rigid structures to promote linguistic diversity and reduce the risk od overfitting to narrow prompt patterns. An example of $S_{in}$ is “Around the campus park, students sit to have lunch, talk with others, and wave at passing friends.” Additional examples are provided in the supplementary material. We use OpenAI’s gpt-4.1 [59] and apply two sequential queries ($Q1$ & $Q2$), each constrained to return JSON-formatted outputs; exact prompts and parameters are listed in the supplementary material.
@@
Generating the Environment ($Q1$). The first query prompts the LLM to generate a plausible environment layout from $S_{in}$, specifying Locations relevant to the scenario. For example, the output may include places like a “coffee shop”, “park”, or “entrance”, together with their category, position, and scale.
@@
Authoring Crowd Events ($Q2$). The second query defines the sequence of agent actions and interactions, both with the environment and with other agents. It takes as input $S_{in}$ together with a simplified list of the previously generated environment, containing only the location names for context. The output is a list of events.
@@
By combining responses from both queries, we construct a high-level crowd scenario $CS$ from the textual input $S_{in}$. We then iterate over the generated events to extract each agent’s actions and group memberships. As LLMs lack precise temporal detail, event sequences are derived from initiation constraints, and action durations are sampled from action-specific distributions. This yields a record $Rec_{i}$ for every agent $i$, capturing its actions and interactions throughout its lifespan in the simulated environment.

### v4

@@ Abstract
Over the past two decades, researchers have made significant steps in simulating agent-based human crowds, yet most efforts remain focused on low-level tasks such as collision avoidance, path following, and flocking. As a result, these approaches often struggle to capture the high-level behaviors that emerge from sustained agent-agent and agent-environment interactions over time. We introduce Generative Crowds (Gen-C), a generative framework that produces crowd scenarios capturing agent-agent and agent-environment interactions, shaping coherent high-level crowd plans. To avoid the labor-intensive process of collecting and annotating real crowd video data, we leverage Large Language Models (LLMs) to bootstrap synthetic datasets of crowd scenarios. To represent those scenarios, we propose a time-expanded graph structure encoding actions, interactions, and spatial context. Gen-C employs a dual Variational Graph Autoencoder (VGAE) architecture that jointly learns connectivity patterns and node features conditioned on textual and structural signals, overcoming the limitations of direct LLM generation to enable scalable, environment-aware multi-agent crowd simulations. We demonstrate the effectiveness of our framework on scenarios with diverse behaviors such as a University Campus and a Train Station, showing that it generates heterogeneous crowds, coherent interactions, and high-level decision-making patterns consistent with the provided context.
@@ 1 Introduction
Simulating human crowds has long been an active research field, producing numerous models and tools for character behavior generation [46, 29]. Several commercial solutions integrated into game engines and rendering software enable large-scale crowd rendering [10, 37, 61, 8] in virtual scenes. While these approaches have advanced visual plausibility, they primarily target motion synthesis rather than the decision processes and interactions that govern behaviors over time.
Most agent-based methods focus on low-level navigation tasks such as collision avoidance, path following, and group steering [60, 38]. While effective for local interactions, they struggle to model higher-level behaviors that arise as individuals pursue goals, interact socially, or respond to environmental cues. Examples include stopping to chat, browsing shop windows, or queuing for a train, activities that require planning and coordination but are rarely represented in current systems, leading to repetitive and less adaptive dynamics. Recent advances such as Meta’s WorldGen [62] and Google’s Genie [3] highlight the need for scalable virtual worlds with coordinated agents.
@@
In this work, we introduce Generative Crowds (Gen-C), a framework for synthesizing high-level behaviors of virtual agents for diverse multi-agent crowd simulations. Rather than targeting low-level physical fidelity, this work emphasizes the generation of plausible, diverse, and goal-driven high-level agent behaviors for autonomously populating complex environments. At the core of our approach lies a crowd scenario graph: a time-expanded graph that encodes the spatial and temporal evolution of the crowd, capturing agent-agent and agent-environment interactions. With the advent of data-driven methods, an obvious approach is to learn such graphs directly from real data. Several works have modeled crowds by leveraging trajectories extracted from limited real-world datasets [32, 26, 33, 5, 42]. However, the cost and effort required to collect and annotate such data, together with its limited coverage of high-level behaviors, quickly become a bottleneck and restrict generalization.
To reduce this dependency, we use Large Language Models (LLMs) [2] to automatically generate a small seed set of representative crowd scenarios. Although LLMs could directly script crowd behaviors, this approach scales poorly, demands extensive prompt design, and lacks efficiency for structured generation. Instead, we use LLMs only to bootstrap initial scenarios, represented as crowd scenario graphs. A generative model then learns a distribution of crowd actions and interactions from these graphs to synthesize new ones. Since existing graph embedding algorithms target static graphs and fail to model dynamic crowd scenarios, we introduce an architecture with two synergistic Variational Graph Autoencoders (VGAEs) [23] that jointly capture agent-agent and agent-environment dynamics. The trained model can be conditioned on natural language, enabling virtual crowd generation directly from text.
Figure 1: Framework Overview. We use a Large Language Model to generate synthetic datasets of crowd scenarios, which we represent as crowd scenario graphs to train a generative model. Conditioned on textual inputs, our model samples these graph representations to populate virtual worlds with diverse crowd scenarios.
@@
Our approach builds upon recent advances in generative AI, where models have been successfully applied to text-driven human motion synthesis [72] and agent or robot planning [36]. However, these methods mainly focus on single-agent scenarios [64, 59, 67, 18, 63]. Closer to our work, text-based diffusion models have been explored for coordinating groups [17] and synthesizing crowd trajectories [50], but they remain limited to unstructured motion patterns and a narrow range of behaviors. In contrast, Gen-C targets multi-agent settings and emphasizes the generation of diverse, high-level behavior compositions and interactions conditioned on scene context, enabling richer and more expressive characters in virtual scenes. Overall, we introduce: (i) a graph-based representation that captures temporal agent-agent and agent-environment interactions for virtual crowds; (ii) a text-conditioned dual-VGAE architecture that jointly learns graph structure and features, enabling scalable generation of novel multi-agent crowd scenarios; and (iii) a synthetic data pipeline, leveraging LLMs to bootstrap crowd scenarios, reducing reliance on real-world annotations.
@@ 2 Related Work
Crowd Simulation. The study of simulating crowds has a long history, with early works introducing rule-based models for navigation and flocking [52], or pedestrian activities [54], while more complex character behaviors, such as those in computer games, are often authored using state machines [66]. Subsequent research proposed higher-level control mechanisms, such as crowd patches and sketch-based navigation fields [68, 19, 45], which allowed users to shape collective motion in structured environments. A major thrust of later work has been to increase plausibility by modeling heterogeneity, for example through group-level properties [51], individual personality traits [7, 21, 12], or stylistic variations across different types of crowds [20]. These methods highlighted the importance of diversity for believable simulations but typically required manual specification or extensive parameter tuning, motivating stochastic authoring approaches that provide direct user control over high-level crowd activity distributions [39]. In parallel, data-driven approaches emerged that learn behavioral patterns from real-world crowd trajectories [25, 31, 71, 30], sometimes adopting graph-based representations [24, 4] that connect crowd modeling to broader developments in machine learning. More recently, reinforcement learning and hybrid imitation-learning methods [43, 42, 50, 5], diffusion-based generative approaches [16, 1], and graph network-based methods [55] have demonstrated the potential of scalable and adaptive policies that capture richer dynamics. Despite advances in crowd simulation, most existing methods focus on low-level navigation or local diversity, neglecting structured high-level multi-agent behaviors driven by agent-agent and agent-environment interactions. Our framework automatically generates heterogeneous, scenario-driven crowd behaviors capturing these multi-layered interactions.
@@
Large Language Models and Autonomous Agents. LLMs have been shown to be effective in domains beyond Natural Language Processing (NLP) such as task planning, embodied reasoning, autonomous agents, and floor plan design [15, 57, 48, 58, 44, 73]. Here, we explore an architecture that allows an LLM to generate a sequence of crowd behaviors, over time and space, that forms synthetic crowd datasets. We further exploit the power of generative NLP models to enable text-guided synthesis of high-level crowd behaviors. Along these lines a lot of work has explored generative models for text-driven human motion generation tasks [72]. Despite impressive results, including generating scene-and-language conditioned motions [64, 59, 67, 18, 63], such methods typically focus on individual characters. But a human crowd is much more than that, consisting of multiple agents performing a sequence of actions while interacting with other agents and the environment. Along these lines, the recent work of Ji et al. [17] uses a text-based diffusion model combined with LLMs to steer groups of agents to predefined destinations, while Lemonari et al. [28] proposes a learning paradigm for mapping text to steering behavior parameters. Our work seeks to learn high-level behaviors captured via crowd scenario graphs that can complement the low-level steering capabilities of such agents.
@@
Graph Generation. Beyond our domain, learning graph representations is fundamental to many real-world applications [65]. While much prior work has focused on static graphs using models like VGAEs [23], recent efforts extend to dynamic graphs whose structures evolve over time. These include approaches using neural temporal point processes, attention mechanisms [70], and deep architectures such as Temporal Graph Networks [53] and Spatio-Temporal Graph Convolutional Networks [69] for modeling temporal dependencies and forecasting. While related, our work focuses on crowd scenario graphs: dynamic graphs with rich relational features that capture the evolution of crowds. Unlike most dynamic graph models aimed at tasks such as link prediction, node classification, or recommendation, our formulation targets the evolution of crowd interactions. Closely related, conditional graph generation has also gained attention, with most approaches generating only the graph structure while assuming fixed node and edge features. For instance, NGG [9] employs conditioned latent diffusion, GraphVAE [56] integrates conditional codes, and MOOD [27] guides diffusion sampling with property gradients. Sequential models [35] use conditional inputs for nodes and edges initialization, while others embed conditions within message-passing schemes [34]. Complementing these approaches, our framework jointly models graph structure and node features through separate, condition-aware latent distributions, enabling generation directly from text.
@@ 3 Overview
Preliminaries. A crowd scenario graph represents a crowd scenario, consisting of agents, a series of events, environment locations, and the associated actions and interactions. To simplify the generation and learning process, we define a finite set of actions and locations which, when combined, can describe a wide range of crowd behaviors. Specifically, an Action is a label from $\mathbf{\mathcal{A}ct}=$~{wait, sit, wander, queue, enter/exit, wave at, discuss, meet, object interact, talk on phone, read, look at, carry} that specifies an agent’s behavior at a given time. When multiple agents perform a shared action (e.g., discuss), this is referred to as an Interaction. An Environment Location is defined by its name, category, position, and scale. While the name may be abstract, the category is selected from a finite set: $\mathbf{\mathcal{L}oc}=$~{building, room, entrance, exhibit, furniture, outdoor area, item, service area}. Finally, an Event encodes a sequence of actions together with their location and participating agents. Specifically, each event consists of a set of action-location pairs (e.g., {discuss, room}), the participating agents, and a trigger condition that determines its initiation (i.e., starts after a specific event).
The Gen-C Framework. Our framework for generating crowd scenarios consists of three main components (Figure 1): Synthetic Data Generation, Learning, and Scenario Generation. First, we use an LLM to produce diverse synthetic scenarios through targeted queries that contains events describing agent actions, interactions, and environment placements (Section 4.1). These scenarios are transformed into graph structures encoding action sequences, relationships, and interactions (Section 4.2). Second, we train two synergistic VGAE models: one reconstructs the graph structure (agent interactions), while the other defines node features (agent actions and navigation). A text-conditioned prior integrates natural language, aligning the generated graphs with input descriptions (Section 5). Finally, novel scenarios are synthesized by sampling from the learned latent distributions, producing contextually consistent multi-agent crowd behaviors (Section 6).
@@ 4 Synthetic Data Generation
Simulating diverse high-level crowd behaviors requires large amounts of high-quality data. However, collecting and annotating real-world crowd data is costly and labor-intensive, and existing crowd datasets are limited in scope, focus primarily on trajectory-level information, and lack high-level semantic annotations. To mitigate these limitations, we explore the use of recent advances in LLMs.
@@ 4.1 Bootstrapping Crowd Scenarios via LLMs
To generate simulations that capture representative crowd dynamics, we design two tailored LLM queries, each serving a specific purpose. The process begins with a single-sentence input $S_{in}$ describing a crowd scenario, which serves as the seed for generation. We construct a set of $100$ such seeds and, for each, create $50$ paraphrased variants to increase diversity, yielding $5000$ textual inputs in total per theme; in this work we introduce two themes, University Campus and Train Station. Each sentence encodes three main components: a location phrase, one or more subject noun phrases, and one or more action verb phrases, without following a fixed syntactic template. We avoid rigid structures to promote linguistic diversity and reduce the risk of overfitting to narrow prompt patterns. An example of $S_{in}$ is “Around the campus park, students sit to have lunch, talk with others, and wave at passing friends.” Additional examples, along with exact prompts and parameters, are provided in the supplementary material. We use OpenAI’s gpt-4.1 [41] and apply two sequential queries ($Q1$ and $Q2$), each constrained to return JSON-formatted outputs.
@@
Generating the Environment ($Q1$). The first query prompts the LLM to generate a plausible environment layout from $S_{in}$, specifying Locations relevant to the scenario. The output may include places like a “coffee shop”, “park”, or “entrance”, together with their category, position, and scale.
@@
Authoring Crowd Events ($Q2$). The second query defines the sequence of agent actions and interactions, both with the environment and with other agents. It takes as input $S_{in}$ together with a simplified list of the previously generated environment, containing only the location names for context. The output is a list of events, as previously defined in Section 3.
@@
By combining responses from both queries, we construct a high-level crowd scenario $CS$ from an $S_{in}$. We then iterate over the generated events to extract each agent’s actions and group memberships. As LLMs lack precise temporal detail, event sequences are derived from initiation constraints, and action durations are sampled from action-specific distributions. This yields a record $Rec_{i}$ for every agent $i$, capturing its actions and interactions throughout its lifespan in the scene.
Figure 2: Per-agent action transitions from LLM generations and real-world annotations, and deviation between the two ($\Delta$) in Train Station theme. Rows represent the “from” action where columns the “to”.
Assessment of LLM-bootstrapped Data. After using an LLM to generate crowd scenarios, we evaluate the reliability and fidelity of the results by annotating high-level action plans of individuals interacting in real train station environments, using a curated set of YouTube videos totaling approximately 30 minutes of footage. Comparable data for university campus environments are scarce, primarily due to ethical restrictions on video collection, and existing datasets such as UCY Dataset [31] exhibit limited diversity in observed actions. Specifically, we divide each video clip into $5$-second windows and assign an action label from $\mathbf{\mathcal{A}ct}$ to each person visible in the frame.
[…]
As an additional statistic, we measure diversity of action plans using the normalized entropy of per-agent action sequences. Let $\mathcal{S}$ be the set of distinct action sequences and $p(s)$ their empirical frequencies. We compute $H(p)\;=\;-\sum_{s\in\mathcal{S}}p(s)\,\log p(s)$ and report the normalized value $\hat{H}=H(p)/\log|\mathcal{S}|\in[0,1]$: $\hat{H}=0$ means one pattern dominates; $\hat{H}=1$ means all patterns are equally likely. By this measure, LLM sequences closely match the annotated real-world data: $\hat{H}_{\text{LLM}}=0.949$ $(95\%\ \text{CI }[0.916,\,0.970])$ vs. $\hat{H}_{\text{Real}}=0.936$ $(95\%\ \text{CI }[0.935,\,0.939])$. The overlapping intervals indicate that LLM data follow real-world sequence variety for the given context. Finally, it is worth noting that the reliability of synthetic LLM data should be interpreted at Gen-C’s level of abstraction: semantically plausible, context-aware high-level behavioral plans, rather than distributional equivalence to real-world trajectories. Accordingly, this evaluation should be viewed as a sanity check, although it is limited by its short duration and the partly subjective mapping to discrete action categories. More systematic real-world grounding through extensive evaluation is needed to fully establish the realism and naturalness of LLM-generated scenarios, which we identify as an important direction for future work.
Figure 3: Snapshot of a crowd scenario graph. A node encodes an action, a location, and the agent ID along with a timestep as i_ti\_t.
Figure 4: Canonical node ordering (top) vs. arbitrary node ordering (bottom).